Circular No. 6183

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

GRO J1629–49 R. H. McNaught, Anglo-Australian Observatory, reports that he and

G. J. Garradd obtained unfiltered CCD images around July 6.6 UT with the 1.0-m telescope at Siding Spring, covering 1.2 deg^{2} around the EGRET position reported on IAUC 6182. Examination of the close vicinity of PMN J1628–4936 indicates the presence of a new source of mag ∼ 18.0 at α = 16^{h}28^{m}57^{s}.86, δ = −49^{o}36′53′′.4 (equinox 2000.0; uncertainty 0′′.1 in each coordinate, based on 14 GSC stars). A star of mag ∼ 19.5 is offset from the suspect by ∼ 2′′ in p.a. 300^{o} and a star of mag ∼ 17.5 has end figures 57^{s}.79, 58′′.9. No image appears in this position on the ESO R or B or the SERC J surveys.

Further to IAUC 6182, G. Kanbach adds that the source was detected in observations made during 1995 June 24–July 2.

Corrigendum. On IAUC 6182, line 8, for Parkes-Molonglo-NRAO read Parkes-MIT-NRAO.

4C 38.41 S. Bosio, G. De Francesco, G. Ghisellini, G. Latini, C. M. Raiteri,

and M. Villata, Torino Observatory, report: “The γ-ray blazar 4C 38.41 (1633+382) underwent a major optical flare, reaching its historical maximum. Magnitudes in the R (Cousins) and B (Johnson) bands obtained with the Pino Torinese 1.05-m telescope (+ CCD camera) are: May 24.02 UT, R = 17.18 ± 0.02; June 6.01, 16.93 ± 0.02; 19.01, 16.70 ± 0.04; 20.95, 16.91 ± 0.04; 27.95, 14.96 ± 0.02; 28.00, 14.96 ± 0.02; 28.88, 16.04 ± 0.02; 29.92, 16.26 ± 0.02; July 4.98, 16.26 ± 0.02; June 28.91, B = 16.98 ± 0.04; 29.89, 17.13 ± 0.03. Follow-up observations at other (especially radio) bands are encouraged.”

TIME ADJUSTMENT ON 1995 DECEMBER 31 Bulletin C10 of the International Earth Rotation Service announces

that a positive leap second will be introduced such that the sequence of UTC second markers will be: 1995 December 31^{d}23^{h}59^{m}59^{s}, 31^{d}23^{h}59^{m}60^{s}, 1996 January 1^{d}00^{h}00^{m}00^{s}. Beginning 1996 January 1, the difference UTC–TAI = −30^{s}.

1995 July 6 Daniel W. E. Green